e3:

The cost in dollars, y, of a large pizza with x toppings from Pat’s Pizzeria can be modeled by a linear function. A large pizza with no toppings costs $14.00. A large pizza with 2 toppings costs $17.50. What is the cost of a pizza with 5 toppings? Round to the nearest penny.

justus:

First of all, Yummy you made me hungry We can turn the question into data points: (0, 14) and (2,17.50). We should solve the two points slope, which is (17.50-14)/2 equal to 1.75. Therefore replacing the standard formula in the point-sope form: y-y1 = m (x-x1) and y-14 = 1.75 x, if x is 5, y is equal to $22.75

e3:

Thank you. Your are so helpful

justus:

No problem! Its a pleasure